# SUREL+: Moving from Walks to Sets for Scalable Subgraph-based Graph Representation Learning

### 17 Dec 2023 v2

@@ Abstract.
Subgraph-based graph representation learning (SGRL) has recently emerged as a powerful tool in many prediction tasks on graphs due to its advantages in model expressiveness and generalization ability. Most previous SGRL models face computational issues associated with the high cost of subgraph extraction for each training or test query. Recently, SUREL was proposed to accelerate SGRL, which samples random walks offline and joins these walks online as a proxy of subgraphs for representation learning. Thanks to the reusability of sampled walks across different queries, SUREL achieves state-of-the-art performance in terms of scalability and prediction accuracy. However, SUREL still suffers from high computational overhead caused by node redundancy in sampled walks. In this work, we propose a novel framework SUREL+ that upgrades SUREL by using node sets instead of walks to represent subgraphs. This set-based representation avoids repeated nodes by definition, but node sets can be irregular in size. To address this issue, we design a customized sparse data structure to efficiently store and index node sets, and provide a specialized operator to join them in parallel batches. SUREL+ is modularized to support multiple types of set samplers, structural features, and neural encoders to complement the structure information loss after the reduction from walks to sets. Extensive experiments have been performed to validate SUREL+ in the prediction tasks of links, relation types, and higher-order patterns. SUREL+ achieves 3-11$\times$ speedups of SUREL while maintaining comparable or even better prediction performance; compared to other SGRL baselines, SUREL+ achieves $\sim$20$\times$ speedups and significantly improves the prediction accuracy.
@@
PVLDB, 16(11): 2939-2948, 2023.
@@ 1. Introduction
Graphs are widely used to model interactions in natural sciences and relationships in social life (Koller et al., 2007; Jumper et al., 2021). Graph-structured data in the real world are highly irregular and often large-scale. To solve inference tasks on graphs, graph representation learning (GRL) that studies quantitative representations of graph-structured data has attracted much attention (Hamilton et al., 2017b; Hamilton, 2020; Wu et al., 2022). Recently, subgraph-based GRL (SGRL) has become an important research direction for researchers studying GRL algorithms and systems, as it achieves far better prediction performance than other approaches in many GRL tasks, especially those involving a set of nodes. Given a set of nodes of interest, namely a queried node set, SGRL models such as SEAL (Zhang and Chen, 2018; Zhang et al., 2021), GraIL (Teru et al., 2020), and SubGNN (Alsentzer et al., 2020) first extract a subgraph around the queried node set (called query-induced subgraph), and then use neural networks to encode extracted subgraphs for prediction. Extensive work shows that SGRL models are more robust (Zeng et al., 2021) and more expressive (Bouritsas et al., 2022; Frasca et al., 2022); while canonical graph neural networks (GNNs) including GCN (Kipf and Welling, 2017) and GraphSAGE (Hamilton et al., 2017a) usually fail to make accurate predictions, due to their limited expressive power (Zhang et al., 2021; Garg et al., 2020; Chen et al., 2020), incapability of capturing intra-node distance information (Srinivasan and Ribeiro, 2020; Li et al., 2020), and improper entanglement between receptive field size and model depth (Huang and Zitnik, 2020; Zeng et al., 2021; Yin et al., 2022). An example in Fig. 2 illustrates how SGRL works for link prediction and demonstrates its advantages over GNNs. Here, canonical GNNs generate and aggregate node-wise representations to predict links, which would map structurally symmetric nodes into the same representation and lead to the ambiguity issue (Xu et al., 2019; Zhang et al., 2021). So far, the advantages of SGRL methods have been proved in many applications, such as link and relation prediction (Zhang and Chen, 2018; Zhang et al., 2021; Teru et al., 2020), higher-order pattern prediction (Meng et al., 2018; Liu et al., 2022), temporal network modeling (Wang et al., 2021), recommender systems (Zhang and Chen, 2020), anomaly detection (Alsentzer et al., 2020; Cai et al., 2021), graph meta-learning (Huang and Zitnik, 2020), subgraph matching (Liu et al., 2020; Lou et al., 2020), and molecular/protein research in life sciences (Wang and Zhang, 2021).
@@
Figure 1. GNNs cannot correctly predict whether $x$ is more likely linked with $y$ or $z$: without distinct features, $y$ and $z$ have the same node representations. The representations based on one-hop neighbors (query-induced subgraph) are more expressive in distinguishing these two node pairs.
@@
Figure 1. GNNs cannot correctly predict whether $x$ is more likely linked with $y$ or $z$: without distinct features, $y$ and $z$ have the same node representations. The representations based on one-hop neighbors (query-induced subgraph) are more expressive in distinguishing these two node pairs.
@@
Figure 2. Overview of SUREL+: a side-by-side comparison with previous SGRL models. For a query, SGRL first prepares the input subgraph paired with structural features, which are then fed into a neural encoder to obtain embeddings for prediction. SUREL+ samples node sets as input, while SEAL (Zhang and Chen, 2018; Zhang et al., 2021) extracts induced subgraphs and SUREL (Yin et al., 2022) samples walks. To compensate for structure loss by reducing subgraphs to node sets, SUREL+ provides various types of set samplers, structure encoders, and set neural encoders. To better serve set-based representations, SUREL+ designs a customized sparse data structure SpG to store sampled node sets with fast access, and supports their online joining in parallel via a sparse join operator SpJoin.
@@
Albeit with multiple benefits of its algorithm, SGRL methods currently face two major computational challenges: (1) Query Dependency. A subgraph must be extracted for each queried node set, which is not reusable across different queries, and cannot be preprocessed if the query is unknown; (2) Irregularity. The extracted subgraphs are irregularly sized, resulting in poor batch processing and load-balancing performance. As shown in Fig. 3 (a), subgraph extraction in SEAL (Zhang and Chen, 2018; Zhang et al., 2021) is prohibitively slow to be deployed in practice. This inspired recent work on dedicated hardware acceleration for subgraph extraction (DGL, 2022; PyG, 2022). However, how to fundamentally improve the scalability and efficiency of SGRL methods remains largely unexplored.
@@
Figure 3. (a) Subgraph extraction in SEAL (Zhang and Chen, 2018; Zhang et al., 2021) has much higher complexity than other simplified form samplers. The size of induced subgraphs is irregular, growing exponentially with the number of extracted hops (Chen et al., 2018; Zeng et al., 2021). (b) Breaking subgraphs into reusable walks reduces complexity but faces the issue of heavy node duplicates.
@@
SUREL (Yin et al., 2022) is the state-of-the-art (SOTA) framework that applies algorithm and system co-design to implement SGRL. It employs the join of node-level sampled walks to represent query-specific subgraphs. Specifically, SUREL treats each node as a seed and runs multiple random walks from the seed on the graph offline. Given a queried node set, SUREL online joins and encodes the sampled walks for all queried nodes for prediction. The join operation builds the connection between sampled walks of queried nodes so that their joint walks can function as the query-induced subgraph. To compensate for the structure loss by representing subgraphs in walks, a structural feature termed relative position encoding (RPE) is adopted to record the position of each distinct node in sampled walks relative to the seed. The RPE is pre-computed offline and attached to sampled walks before being fed into neural networks (NNs) to make predictions. In SUREL, sampled walks from one seed can be reused for multiple queries whenever that seed node is involved. SUREL significantly improves the scalability of SGRL by this walk-sharing mechanism and benefits from the regularity of walks, enabling highly parallel walk sampling and online joining through a dedicated system design. However, SUREL still faces several inherent drawbacks of the walk-based representation, namely high node redundancy in sampled walks (over 55% duplication, see Fig. 3 (b)). This further raises the following issues: (1) extra space for hosting walks and positional encoding in memory, (2) extra time for operations on repeated nodes in subsequent routines of walk joining and NN-based encoding, and (3) high workload of data transfer between CPU and GPU.
@@
In this work, we upgrade SUREL and develop a novel framework SUREL+ that again benefits from algorithm-system co-design. The core concept of SUREL+ is simple: instead of using walks to represent subgraphs, we now employ node sets, thereby obviating node duplication. However, this new idea based on node sets introduces several algorithm and system design difficulties. From an algorithmic perspective, as explored in this study, the transition from induced subgraphs to walks and then to node sets results in a considerable loss of structural information. Therefore, the first priority is to develop a method that can compensate for structural losses while maintaining performance. On the system side, SUREL (Yin et al., 2022) utilizes walks, which can be easily stored and processed in an aligned format by controlling the length and number of sampled walks. In contrast, node sets are irregular in size, creating difficulties for efficient storage and fast access. To sum up, how to coordinate the designs of both sides constitutes the primary challenge of this work.
@@
SUREL+ tackles the above challenges throughout its entire pipeline. To avoid node duplication in walk sampling, SUREL+ only keeps unique nodes from each seed node’s sampled neighborhood during preprocessing. To compensate for structure information loss, SUREL+ incorporates various types of set samplers and structure encoders to preserve local structures: set samplers employ different graph metrics to measure node importance and determine sampling rules; structure encoders support landing probabilities of random walks (Li et al., 2019), shortest path distances, and personalized PageRank scores (Jeh and Widom, 2003), covering most of the structural features used by previous SGRL models (Zhang and Chen, 2018; Zhang et al., 2021; Li et al., 2020; Teru et al., 2020; Yin et al., 2022). Furthermore, SUREL+ designs a customized sparse data structure, namely SpG, which can efficiently store sampled node sets and achieve fast access. During training and inference, SUREL+ implements a sparse operator SpJoin to perform join operations on the sampled node sets and their associated structural features of all queried nodes, acting as the query-induced subgraph for prediction. In addition, to capture diverse levels of interactions between node and structural features, SUREL+ supports multiple set neural encoders, such as multi-linear perception (MLP) with mean pooling, set attention (Veličković et al., 2018) and LSTM (Hamilton et al., 2017a) that ensure sufficient expressive power and consistent performance across various types of SGRL tasks.
@@
  • Algorithm: SUREL+ is a novel SGRL framework (open source), that utilizes reusable node sets and associates them with various types of structural features to represent query-induced subgraphs via online joining. Adopting sets in SUREL+ greatly reduces memory and computation costs without degrading prediction performance compared with the SOTA baselines.
  • System: SUREL+ designs a customized sparse data structure SpG and a sparse join operator SpJoin to support efficient storage and processing of node sets, which achieves much better efficiency and scalability than previous SGRL methods.
@@ 2.1. Notations and Relevant Definitions in SGRL
Given a query $Q$ of node set over graph $\mathcal{G}$, SGRL aims to learn a representation of the query-induced subgraph $\mathcal{G}_{Q}$ to make prediction $f(\mathcal{G}_{Q})$. ()($\cdot$)( ⋅ ) is usually a neural network. SGRL tasks come with some labeled queries $\{(Q_{i},y_{i})\}_{i=1}^{L}$ for supervision (positive samples) and other unlabeled queries $\{Q_{i}\}_{i=L+1}^{L+N}$ for inference.
@@
Review of SGRL Methods The current SGRL pipeline mainly has three parts, as shown in the Algorithm Design section of Fig. 2: subgraph preparation, structural feature construction, and neural encoder to obtain subgraph embeddings. Classical SGRL models often group query-dependent parts together (e.g., SEAL (Zhang and Chen, 2018; Zhang et al., 2021) couples subgraph extraction with labeling trick (Zhang et al., 2021)), and then apply neural encoder on extracted subgraphs with attached structural features for prediction. However, such coupling is expensive and makes the computed intermediate results not reusable across queries, which motivates recent SGRL methods to decouple them. SUREL (Yin et al., 2022) substitutes explicit subgraph extraction with the online joining of multiple walks pre-sampled offline associated with position encoding defined on sampled walks as structural features. Sampled walks and positional encoding are node-level and can be reused to serve multiple queries, which greatly improves efficiency. Lastly, neural networks are applied to encode and aggregate the embedding of joint walks for prediction.
@@ 2.2. Related Works
Scalable SGRL Design. Recent works on SGRL models have primarily focused on efficient subgraph extraction. Various techniques have been proposed, including PPR-based (Bojchevski et al., 2020; Zeng et al., 2021) and random walk-based (Yin et al., 2022) subgraph samplers, node neighborhood sampling through CUDA kernel (DGL,  (DGL, 2022)) and tensor operations (PyG, (PyG, 2022)). Some frameworks have customized data structures to better support subgraph operations and gain higher throughput, such as associative arrays in SUREL (Yin et al., 2022), temporal-CSR in TGL (Zhou et al., 2022), and GPU-orientated dictionary in NAT (Luo and Li, 2022). To achieve scalable modeling design, GDGNN (Kong et al., 2022) utilizes node representations along the geodesic path between queried nodes for prediction, partially decoupling structural feature construction from subgraph extraction. BUDDY (Chamberlain et al., 2023) employs subgraph sketches to avoid explicitly constructing subgraphs for link prediction. However, these works either focus on specific aspects of the SGRL scalability issue, namely extraction, storage, and feature construction bottlenecks, or are limited to one type of SGRL tasks. In contrast, SUREL+ provides a comprehensive co-design approach in scalable sampling, efficient storage, and expressive modeling, offering a general subgraph-based framework for scalable SGRL.

### 27 Dec 2023 v3

@@ Abstract.
Subgraph-based graph representation learning (SGRL) has recently emerged as a powerful tool in many prediction tasks on graphs due to its advantages in model expressiveness and generalization ability. Most previous SGRL models face computational challenges associated with the high cost of subgraph extraction for each training or test query. Recently, SUREL was proposed to accelerate SGRL, which samples random walks offline and joins these walks online as a proxy of subgraph for representation learning. Thanks to the reusability of sampled walks across different queries, SUREL achieves state-of-the-art performance in terms of scalability and prediction accuracy. However, SUREL still suffers from high computational overhead caused by node duplication in sampled walks. In this work, we propose a novel framework SUREL+ that upgrades SUREL by using node sets instead of walks to represent subgraphs. This set-based representation eliminates repeated nodes by definition but can also be irregular in size. To address this issue, we design a customized sparse data structure to efficiently store and access node sets and provide a specialized operator to join them in parallel batches. SUREL+ is modularized to support multiple types of set samplers, structural features, and neural encoders to complement the structural information loss after the reduction from walks to sets. Extensive experiments have been performed to validate SUREL+ in the prediction tasks of links, relation types, and higher-order patterns. SUREL+ achieves 3-11$\times$ speedups of SUREL while maintaining comparable or even better prediction performance; compared to other SGRL baselines, SUREL+ achieves $\sim$20$\times$ speedups and significantly improves the prediction accuracy.
@@
Haoteng Yin, Muhan Zhang, Jianguo Wang, Pan Li. PVLDB, 16(11): 2939-2948, 2023.
@@ 1. Introduction
Graphs are widely used to model interactions in natural sciences and relationships in social life (Koller et al., 2007; Jumper et al., 2021). Graph-structured data in the real world are highly irregular and often large-scale. To solve inference tasks on graphs, graph representation learning (GRL) that studies quantitative representations of graph-structured data has attracted much attention (Hamilton et al., 2017b; Hamilton, 2020; Wu et al., 2022). Recently, subgraph-based GRL (SGRL) has become an important research direction for researchers studying GRL algorithms and systems, as it achieves far better prediction performance than other approaches on many GRL tasks, especially those involving a set of nodes. Given a set of queried nodes, SGRL models such as SEAL (Zhang and Chen, 2018; Zhang et al., 2021), GraIL (Teru et al., 2020), and SubGNN (Alsentzer et al., 2020) first extract a subgraph around the queried node set (called query-induced subgraph), and then use neural networks to encode extracted subgraphs for prediction. Extensive work shows that SGRL models are more robust (Zeng et al., 2021) and more expressive (Bouritsas et al., 2022; Frasca et al., 2022); while canonical graph neural networks (GNNs) including GCN (Kipf and Welling, 2017) and GraphSAGE (Hamilton et al., 2017a) usually fail to make accurate predictions, due to their limited expressive power (Zhang et al., 2021; Garg et al., 2020; Chen et al., 2020), incapability of capturing intra-node distance information (Srinivasan and Ribeiro, 2020; Li et al., 2020), and improper entanglement between receptive field size and model depth (Huang and Zitnik, 2020; Zeng et al., 2021; Yin et al., 2022). An example in Fig. 2 illustrates how SGRL works for link prediction and demonstrates its advantages over GNNs. Here, canonical GNNs generate and aggregate node-wise representations to predict links, which would map structurally symmetric nodes without distinct features into the same representation and thus lead to the ambiguity issue (Xu et al., 2019; Zhang et al., 2021). So far, the advantages of SGRL methods have been verified in many applications, such as link and relation prediction (Zhang and Chen, 2018; Zhang et al., 2021; Teru et al., 2020), higher-order pattern prediction (Meng et al., 2018; Liu et al., 2022), temporal network modeling (Wang et al., 2021), recommender systems (Zhang and Chen, 2020), anomaly detection (Alsentzer et al., 2020; Cai et al., 2021), graph meta-learning (Huang and Zitnik, 2020), subgraph matching (Liu et al., 2020; Lou et al., 2020), and molecular/protein study in life sciences (Wang and Zhang, 2021).
@@
Figure 1. GNNs cannot correctly predict whether $x$ is more likely linked with $y$ or $z$: $y$ and $z$ have the same representation without distinct features. The representations based on one-hop neighbors (query-induced subgraph) are more expressive in distinguishing node pairs $(x,y)$ and $(x,z)$.
@@
Figure 1. GNNs cannot correctly predict whether $x$ is more likely linked with $y$ or $z$: $y$ and $z$ have the same representation without distinct features. The representations based on one-hop neighbors (query-induced subgraph) are more expressive in distinguishing node pairs $(x,y)$ and $(x,z)$.
@@
Figure 2. Overview of SUREL+: a side-by-side comparison with previous SGRL models. For a query, SGRL first prepares the input subgraph attached with structural features, which are fed into a neural encoder to obtain embeddings for prediction. SUREL+ samples node sets as input, while SEAL (Zhang and Chen, 2018; Zhang et al., 2021) extracts query-induced subgraphs and SUREL (Yin et al., 2022) runs random walks. To compensate for the structure loss by reducing subgraphs to node sets, SUREL+ provides various types of set samplers, structure encoders, and set neural encoders. To better serve set-based representations, SUREL+ designs a customized sparse data structure SpG to store sampled node sets with fast access, and supports their online joining in parallel via a sparse join operator SpJoin.
@@
Albeit with multiple benefits of its algorithm, SGRL methods currently face two major computational challenges: (1) Query Dependency. A subgraph must be extracted for each queried node set, which is not reusable across different queries, and cannot be preprocessed if the query is unknown; (2) Irregularity. The extracted subgraphs are irregularly sized, resulting in poor batch processing and load-balancing performance. As Fig. 3 (a) shows, subgraph extraction in SEAL (Zhang and Chen, 2018; Zhang et al., 2021) is prohibitively slow for practical deployment. This inspired recent work on dedicated hardware acceleration for extracting subgraphs (DGL, 2022; PyG, 2022). However, how to fundamentally improve the scalability and efficiency of SGRL methods remains largely unexplored.
@@
Figure 3. (a) Subgraph extraction in SEAL (Zhang and Chen, 2018; Zhang et al., 2021) has much higher complexity than samplers of other simplified forms. The size of query-induced subgraphs is irregular, growing exponentially w.r.t. the number of extracted hops (Chen et al., 2018; Zeng et al., 2021). (b) Breaking subgraphs into reusable walks faces a serious node duplication issue.
@@
SUREL (Yin et al., 2022) is the state-of-the-art (SOTA) framework that applies algorithm and system co-design to implement SGRL. It decouples the input from specific queries by sampling node-level random walks and uses the joint walks of queried nodes as a proxy of subgraphs. Specifically, SUREL treats each node in the graph as a seed and runs multiple random walks from the seed offline. Given a queried node set, SUREL online joins and encodes the sampled walks of all queried nodes for prediction. The join operation builds the connection between sampled walks of queried nodes so that their joins can function as the query-induced subgraph. To compensate for the structure loss by representing subgraphs in walks, a structural feature termed relative position encoding (RPE) is adopted to record the relative distance between each distinct node in sampled walks and the seed. The RPE is pre-computed offline and attached to sampled walks before being fed into neural networks (NNs) to make predictions. Sampled walks from one seed can be reused for multiple queries whenever that seed node is involved. Through this walk-sharing mechanism, SUREL significantly improves the efficiency of SGRL. The regularity of walks also enables highly parallel walk sampling and online joining with a dedicated system design. However, SUREL still faces several inherent drawbacks of the walk-based representation, namely high node redundancy in sampled walks (over 55% is duplicated, see Fig. 3 (b)). This further raises the following issues: (1) extra space for hosting walks and positional encoding in memory, (2) extra time for operations on duplicated nodes in subsequent routines of walk joining and NN-based encoding, and (3) high workload of data transfer between CPU and GPU.
@@
In this work, we upgrade SUREL and develop a novel framework SUREL+ that again benefits from algorithm and system co-design. The core concept of SUREL+ is simple: instead of using walks to represent subgraphs, we now employ node sets, thereby obviating node duplication. However, this new idea based on node sets also introduces several algorithm and system design difficulties. From an algorithmic perspective, as explored in this study, the transition from induced subgraphs to walks and then to node sets results in a considerable loss of structural information. Therefore, the first priority is to develop a method that can compensate for such loss while maintaining performance. On the system side, SUREL (Yin et al., 2022) utilizes walks, which can be easily stored and processed in an aligned format by controlling sampling parameters. In contrast, node sets are irregular in size, creating difficulties for efficient storage and fast access. To sum up, how to coordinate the designs of both sides constitutes the primary challenge of this work.
@@
SUREL+ tackles the above challenges through its entire pipeline. To avoid node duplication in walk sampling, SUREL+ only keeps unique nodes from neighborhood sampling of seed nodes during preprocessing. To compensate for structural information loss, SUREL+ incorporates various types of set samplers and structure encoders to preserve local graph structures: set samplers employ different graph metrics to measure node importance and determine sampling rules; structure encoders support landing probabilities of random walks (Li et al., 2019), shortest path distances, and personalized PageRank scores (Jeh and Widom, 2003), covering most of the structural features used by previous SGRL models (Zhang and Chen, 2018; Zhang et al., 2021; Li et al., 2020; Teru et al., 2020; Yin et al., 2022). Furthermore, SUREL+ designs a customized sparse data structure, namely SpG, which can efficiently store sampled node sets and achieve fast access. A sparse operator SpJoin is developed accordingly to perform join operations on the sampled node sets and associated structural features for serving queries online. To capture diverse levels of interactions between node and structural features, SUREL+ introduces multiple set neural encoders, such as multi-linear perception with mean pooling, set attention (Veličković et al., 2018) and LSTM (Hamilton et al., 2017a) that ensure sufficient expressive power and consistent performance across various types of SGRL tasks.
@@
  • Algorithm: SUREL+ is a novel SGRL framework (open source), which utilizes reusable node sets associated with various structural features to represent query-induced subgraphs via online joining. Compared with the SOTA baselines, the proposed set-based subgraph representation greatly reduces memory and computation costs without degrading prediction performance.
  • System: SUREL+ designs a customized sparse data structure SpG and a sparse join operator SpJoin to support efficient storage and fast access of node sets, which achieves much lower latency and higher throughput than previous SGRL methods.
@@ 2.1. Notations and Relevant Definitions in SGRL
Given a query $Q$ of node set over graph $\mathcal{G}$, SGRL aims to learn a representation of the query-induced subgraph $\mathcal{G}_{Q}$ to make prediction $f(\mathcal{G}_{Q})$. $f(\cdot)$ is usually a neural network. SGRL tasks come with some labeled queries $\{(Q_{i},y_{i})\}_{i=1}^{L}$ for supervision (positive samples) and other unlabeled queries $\{Q_{i}\}_{i=L+1}^{L+N}$ for inference.
@@
Review of SGRL Methods The current SGRL pipeline mainly has three parts, as shown in the Algorithm Design section of Fig. 2: subgraph preparation, structural feature construction, and neural encoder to obtain the readout of subgraphs. Classical SGRL models often group query-dependent parts together (e.g., SEAL (Zhang and Chen, 2018; Zhang et al., 2021) couples subgraph extraction with labeling trick (Zhang et al., 2021)), and then apply GNNs on extracted and labeled subgraphs for prediction. However, such coupling is expensive and makes the computed intermediate results (e.g. labeled subgraphs) not reusable across queries, which motivates recent SGRL methods to decouple them. SUREL (Yin et al., 2022) substitutes explicit subgraph extraction with online joining of multiple pre-sampled walks attached with positional encoding defined on walk landing as structural features, both of which are node-level and thus can be reused to serve multiple queries. Lastly, it applies neural networks to encode joint walks and aggregate their embeddings for prediction.
@@ 2.2. Related Works
Scalable SGRL Design. Recent works on SGRL models have primarily focused on efficient subgraph extraction. Various techniques have been proposed, including PPR-based (Bojchevski et al., 2020; Zeng et al., 2021) and random walk-based (Yin et al., 2022) subgraph samplers, node neighborhood sampling through CUDA kernel (DGL,  (DGL, 2022)) and tensor operations (PyG, (PyG, 2022)). Some frameworks have customized data structures to better support subgraph operations and gain higher throughput, such as associative arrays in SUREL (Yin et al., 2022), temporal-CSR in TGL (Zhou et al., 2022), and GPU-orientated dictionary in NAT (Luo and Li, 2022). To achieve scalable modeling design, GDGNN (Kong et al., 2022) utilizes node representations along the geodesic path between queried nodes for prediction, partially decoupling structural feature construction from subgraph extraction. BUDDY (Chamberlain et al., 2023) employs subgraph sketches to avoid explicitly constructing subgraphs for link prediction. However, these works either focus on specific aspects of computational issues in SGRL, namely bottlenecks of extraction, storage, and feature construction, or are limited to one type of SGRL task. In contrast, SUREL+ provides a comprehensive co-design approach in scalable sampling, efficient storage, and expressive modeling, offering a general and scalable framework for various SGRL tasks.